Zinc will react with hydrochloric acid to give hydrogen in accord with the equation given below. How many mL of a 3M solution HCL are required to react with 10.0 g of zinc? ( Molar mass of Zn=(65.41 g/mol )

Zn(s) + 2 HCL(aq) ====>ZnCl2(aq) + H2 (g)
Answer #1

Zinc moles = mass of Zinc / atomic mass of Zinc = 10 /65.41 = 0.15288

as per reaction 1 Zn reacts with 2HCl.

Hence HCl moles required = 2 x 0.15288 = 0.30576

Now we have Molarity of HCl given = 3M

Molarity = moles of HCl / vol of HCl

3 = 0.30576 /vol of HCl

vol = 0.102 L = 102 ml